OpenStudy (anonymous):

A rectangle has a length of 18 cm and a perimeter of 68 cm. Which equation could be used to find the width of the rectangle? A. w + 18 = 68 B. 18w = 68 C. 2w + 18 = 68 D. 2w + 36 = 68

OpenStudy (anonymous):

i think d

OpenStudy (anonymous):

@roz_mar23

OpenStudy (anonymous):

D because u know 18 is the length. 18*2=36

OpenStudy (anonymous):

Soo 68-36=32/2=16 is the width

OpenStudy (anonymous):

@preteenpony do u get it?